The most important inventory item has annual holding costs of $330.00. One method of shipping this item is 4 days faster than the other, but it is $1.15 more per unit.

How much more expensive is the slower method compared to the faster method? (Include the pennies in your answer)
Answer #1

Per day holding cost = 330/365 = 0.904109589

4 day holding cost = 4*0.904109589 = 3.616438356

So slower method is expensive by (3.616438356-1.15) = 2.466438356 = $2.47 per unit (Rounded to nearest penny)